Let D={(x,y)∈R^2:−5≤x≤5 and 0≤y≤√(25−x^2)}. Then D

is the shape of a
Select one:
a. quarter circle in the second and first quadrant with radius 25.
b. semi-circle in the second and first quadrant with radius 5.
c. semi-circle in the second and first quadrant with radius 25.
d. quarter circle in the second and first quadrant with radius 5.
1
Expert's answer
2020-06-04T19:36:27-0400

Lets consider a function "y= \sqrt{25-x^2}". Then "y^2=25-x^2" and "x^2+y^2=25". It's a circle with the center in (0, 0) and radius 5.

As we have D={(x,y)∈R^2:−5≤x≤5 and 0≤y≤√(25−x^2)}, so our shape is semi-circle in the second and first quadrant with radius 5. The right answer is b.